# Look into the Future: Deep Contextualized Sequential Recommendation

### v1

@@ 4.3. Retriever
The retriever $R$ is responsible for retrieving similar context sequences for target interaction $d_{t}$. We use the feature part $x_{t}$ of $d_{t}$ as the query to utilize the search engine to find similar user contexts in the retrieval pool. This section will cover the detail of the datastore and search processes.
Datastore. The datastore serves as a database in which the keys correspond to all the interactions within the retrieval dataset denoted as $\mathcal{D}_{\text{retrieval}},$ while the values correspond to their associated context sequences. For the sake of expediency, we employ a pretrained encoder to perform offline encoding of all these sequences into their respective embedding representations while the key is still the sparse format for retrieval.
Figure 4. The datastore in LIFT. The primary keys within this datastore are intricately linked to the interactions in the dataset $\mathcal{D}_{\text{retrieval}}$, which represents the original data format. In parallel, the datastore houses values that align with the contextual information associated with these interactions. It is noteworthy that this contextual information is derived through an encoding process that applies a pre-trained encoder, denoted as $E$, to the context sequences.
@@
Query. The retrieval process is to use the target sample’s raw features $x_{t}$ as the query to find the nearby keys in the datastore. We use the traditional BM25 (Robertson et al., 1995) algorithm to retrieve relevant keys, and return top $K$ ranked key-value pairs. In the BM25 algorithm, we treat any sample $x_{z}$ as the query and a key $x_{d}$ in the datastore as the document. As such, the ranking score can be calculated as
@@
Let $\hat{\mathcal{K}}$ be the retrieved keys and $\hat{\mathcal{V}}$ be the retrieved values from the retrieved set. After the retrieval process, we obtain the top-$K$ relevance pairs in the datastore.
@@ 4.4. Predictor
After the retrieval process, we obtain the context of the target sample $x_{t}$. The contextual information includes the user $u_{t}$’s history and both the history and future information of the target sample’s relevant interactions. As shown in Figure 3 right part of stage 2, in the predictor, we use both the contextual information and the embedding of $x_{t}$ to produce the final label prediction $\hat{y_{t}}$.
Key-based Attention Aggregation. From the retriever, we get the context embedding set $\hat{\mathcal{V}}$ of relevant interactions. Before feeding to the final prediction layer, we use the key-based attention mechanism to aggregate the retrieved $\hat{\mathcal{K}}$ and $\hat{\mathcal{V}}$. Let $x_{i}$ be the $i$-th sample in the key set $\hat{\mathcal{K}}$. We use an embedding layer to convert $x_{i}$ into a $M\times w$-dimensional dense vector $\mathbf{x}_{i}$ and we use the same embedding layer to map $x_{t}$ into a $M\times w$-dimensional dense vector $\mathbf{x}_{t}$. The key attention weight is defined as
@@
Prediction. In the final prediction layer, we first use a factor interaction layer to perform the inter-sample interaction for the target sample and the retrieved key set. Since these two vectors are categorical data, one could use high-order interaction further explore the useful patterns, but this is out of the scope of this paper. Finally, we feed the embedding of the target sample, retrieved keys, and retrieved values to an MLP to make the label prediction.
@@ 4.5. Time complexity & Speedup
In this section, we provide a brief time complexity analysis of LIFT framework and then discuss the speedup feasibility.
Time complexity. We aim to provide a comprehensive analysis of the time complexity within our proposed framework, which consists of two major components.
Speedup. To enhance the efficiency of the retrieval process, we may adopt a method of storing the retrieval outcomes, eliminating the need to conduct the retrieval procedure during online inference. This approach shifts the retrieval process to an offline setting. For instance, in recommender systems, this method involves storing the retrieval results for all (or frequent) users, along with their respective recalled candidate items. This method represents a balance between the complexities of space and time, optimizing resource utilization by trading off storage space against processing speed. On the other hand, a recent work (Zheng et al., 2024) has proposed the use of knowledge distillation to distill the datastore of a recommender system into a neural network module, thereby fully avoiding the time and disk space overhead brought by the retrieval module. Thus, it is highly potential that the efficiency of the retrieval process will be sufficiently optimized to make LIFT practically deployable.
@@ 5. Experiments
This section starts with five research questions (RQs), which we use to guide the experiments and discussions. [^1] We provide the experiment code with running instructions on Anonymous GitHub (https://anonymous.4open.science/r/LIFT-277C/).
@@
  • RQ6
    How is the training and inference efficiency of LIFT and whether is it potential to be deployed online?
@@ 5.1. Datasets
We evaluate the performance of LIFT by conducting experiments for CTR prediction tasks on three large-scale real-world datasets, i.e., Taobao, Tmall, and Alipay [^2] https://tianchi.aliyun.com/dataset/x, where ‘x’ is ‘649’, ‘42’, and ‘53’ for Taobao, Tmall, and Alipay, respectively.. We count the number of instances, fields, and users of these datasets, which is shown in Table 2.
Datasets # Instances # Fields # Users
Taobao 100,150,807 4 987,994
Tmall 54,925,331 9 424,170
Alipay 35,179,371 6 498,308
Table 2. Dataset statistics.
Figure 5. Dataset split scheme.
As shown in Figure 5, following the experiment setup of (Qin et al., 2021), we split each dataset into three parts, i.e., retrieval set, train set, and test set, based on the order of the global timestamps. Specifically, the retrieval set consists of the earliest data instances, the test set comprises the latest data instances, and the remaining intermediate data instances are allocated to the train set. As for the pretraining, we use the retrieval set to pretrain the encoder. For the non-retrieval baseline models, we merge the retrieval set and the training set as the final training set.
@@ 5.2. Evaluation Metrics
We choose the widely used metrics area under the ROC curve (AUC) and negative log-likelihood (LogLoss) to evaluate the performance. We also conduct a significance test on the metrics between LIFT and the best performed baseline, marked with “*” for positive results.
@@
Table 5. The influence of the usage of future data in LIFT.
@@ 5.3. Compared Methods
We compare LIFT with nine solid baselines that can be categorized into three groups. The first group is traditional tabular models, which do not utilize the sequential or retrieval mechanism. This group includes GBDT (Chen et al., 1996), a widely-used gradient-boosted trees model, and DeepFM (Guo et al., 2017), a factorization-machine based deep neural network. The second group consists of end-to-end sequential deep models, such as DIN (Zhou et al., 2018b) and DIEN (Zhou et al., 2019) that are attention-based recurrent neural networks for CTR prediction, and BERT4Rec (Sun et al., 2019) that is a Transformers-based sequential recommendation model. The third group includes retrieval-based models, such as SIM (Qi et al., 2020), UBR (Qin et al., 2020), RIM (Qin et al., 2021) and DERT(Zheng et al., 2023). Additionally, FATE (Wu et al., 2021) is a tabular data representation learning model that can be viewed as a random retrieval method for prediction.
Additionally, FATE (Wu et al., 2021) is a tabular data representation learning model, which accepts a random mini-batch of data instances as the entire input and predicts all its elements simultaneously, enabling interactions between the samples in the mini-batch. This approach can be viewed as leveraging a random retrieval method for prediction.
@@ 5.4. Overall Performance (RQ1)
The overall performance comparison result is provided in Table 3, where we have the following observations: (i) LIFT consistently outperforms all the nine baselines. Specifically, compared to the best baseline RIM, LIFT achieves the relatively improved AUC by 3.43%, 1.17% and 9.97% on Taobao, Tmall, and Alipay, respectively. This clearly demonstrates the effectiveness of context information learned and retrieved in LIFT framework. (ii) Without the pretraining, LIFT still consistently outperforms the best baseline, which indicates that the raw information of future and history still yields a significant improvement on the prediction performance. (iii) The retrieval-based methods LIFT and RIM are superior to the traditional methods and sequential CTR models, which means that the retrieval methods are able to make better use of the context information.
@@ 5.5. Further Analysis
Future Context (RQ2). Table 5 compares the different parts of the context information’s impact on the final prediction performance. We can observe that both the history information and future information have a large impact on the final prediction. If we only use part of the context information, such as future only or history only, we will get a worse result than using both of them, which indicates that the future part of the context provides different information compared with the history part. This observation is totally neglected by previous works.
Pretraining (RQ3). From the last two rows of Table 3, we can find that over the three datasets, the model with pretraining performs better than those not, which clearly demonstrates the efficiency of pretraining. The enhanced performance of the pretrained encoder implies that relying solely on the label provided by the target sample is insufficient for effectively capturing contextual information. To improve the modeling of context, it is imperative to harness the inherent information embedded within the context itself.
Hyperparameter in the Retriever (RQ4). In the LIFT framework, $L$ means the context sequence length and $K$ means the retrieved samples count. From Figure 7 and Figure 8, we can find that with the increase of $K$, the final prediction result first goes up and then down. The curve indicates that along with with the $K$, the information first increases, and then more noises are introduced to become the dominant part. The AUC curves of $L$ show the similar trends of $K$, i.e., as $L$ increases, the AUC first gets better and then drops down. We think it is the same reason as for $K$. Moreover, because of the limitation of the GPU resources, we only conduct this hyperparameter study on Taobao and Alipay. and we could only increase the $L$ to 70 in the history length of Taobao, where the downtrend just started.
@@
The Choice of the Encoder (RQ5). The prediction performance of LIFT with different encoders is provided in Table 4, where we can observe that the Transformer decoder based encoder yields the best performance on both non-pretraining and pertraining settings. The reason could be that the Transformer decoder could be aware of the position of the sequence better. This justifies our choice of architecture implementation of the encoder. Further, from Figure 6, we observe the mask ratio’s influence on the prediction performance. Usually, the traditional models that encode the user interaction sequence only use the last label as the supervised signal. From the result, we can find that in all three datasets, such $1$ label supervised strategy is not the best choice. The best choices mask label rate is different over datasets where the number is usually larger than 50%.
Time Efficiency (RQ6). To evaluate the time efficiency of LIFT, we compare the real inference time of it against other three mainstream models in different types. As shown in Figure 9, LIFT does not introduce significant overload compared with the basic retrieval-based model RIM. With offline pertaining, LIFT acts far more efficiently than the transformer-based model BERT4Rec. It is potential to deploy LIFT to the real recommender systems if some speedup schemes are adopted, such as offline retrieval and distillation.
Figure 9. Inference time on Alipay.
Training behavior. The pretraining and training curves of LIFT on three datasets are presented in Figure 10 and 11. In the LIFT model, there are two modules that need to be trained. One is the encoder module used to encode the user sequence. The other is the predictor module that makes prediction by combining user history and information from similar historical and future contexts.
In the LIFT model, our training is divided into two steps: first, we train the encoder module, followed by the predictor module. In Figure 10, we can see that the pre-training of the encoder module converges smoothly across all three datasets. In Figure 11, we observe that the inclusion of the retrieval module makes the neural network’s training curve appear steeper, but it ultimately converges. The curves demonstrate that the pretraining and training processes of LIFT are stable and effective.
Figure 10. Pretraining curves of LIFT.
Figure 11. Training curves of LIFT.
6. Conclusion
In this work, we propose a retrieval-based framework called LIFT to better utilize the context information of the current user interaction. We are the first work to include the future information as the part of the context without data leakage. Moreover, we use a pretraining method to mine the information in the context sequences better and propose a novel mask behavior loss. The performance of the LIFT framework shows that both history and future information yield a significant improvement in the CTR prediction performance. Also, from the comparison of the pretraining method, we find that the context pretraining is a potential way to further improve the prediction performance of a variety of sequential recommendation models. In the future, we will investigate deeper on context representation learning with more sophisticated retrieval methods and the speedup schemes to deploy LIFT to real-world recommenders.

### v2

@@ 4.3. Retriever
The retriever $R$ is responsible for retrieving context sequences similar to the target interaction $d_{t}$. We use the feature part $x_{t}$ of $d_{t}$ as the query to utilize the search engine in finding similar user contexts in the retrieval pool. This section will cover the details of the datastore and search processes.
Datastore. The datastore serves as a database in which the keys correspond to all the interactions within the retrieval dataset denoted as $\mathcal{D}_{\text{retrieval}},$ while the values correspond to their associated context sequences. For the sake of efficiency, we employ a pretrained encoder to perform offline encoding of all these sequences into their respective embedding representations while the key is still the sparse format for retrieval.
@@
Query. The retrieval process uses the target sample’s raw features $x_{t}$ as the query to find the nearby keys in the datastore. We use the traditional BM25 (Robertson et al., 1995) algorithm to retrieve relevant keys, and return the top $K$ ranked key-value pairs. In the BM25 algorithm, we treat any sample $x_{z}$ as the query and a key $x_{d}$ in the datastore as the document. As such, the ranking score can be calculated as
@@
Let $\hat{\mathcal{K}}$ represent the retrieved keys and $\hat{\mathcal{V}}$ be the retrieved values from the retrieved set. After the retrieval process, we obtain the top-$K$ relevance pairs in the datastore.
@@ 4.4. Predictor
After the retrieval process, we obtain the context of the target sample $x_{t}$. The contextual information includes the user $u_{t}$’s history and both the history and future information of the target sample’s relevant interactions. As shown in the bottom part of stage 2 in Figure 3, in the predictor, we use both the contextual information and the embedding of $x_{t}$ to produce the final label prediction $\hat{y_{t}}$.
Key-based Attention Aggregation. From the retriever, we get the context embedding set $\hat{\mathcal{V}}$ of relevant interactions. Before feeding this to the final prediction layer, we use the key-based attention mechanism to aggregate the retrieved $\hat{\mathcal{K}}$ and $\hat{\mathcal{V}}$. Let $x_{i}$ be the $i$-th sample in the key set $\hat{\mathcal{K}}$. We use an embedding layer to convert $x_{i}$ into a $M\times w$-dimensional dense vector $\mathbf{x}_{i}$ and we use the same embedding layer to map $x_{t}$ into a $M\times w$-dimensional dense vector $\mathbf{x}_{t}$. The key attention weight is defined as
@@
Prediction. In the final prediction layer, we first use a factor interaction layer to perform the inter-sample interaction for the target sample and the retrieved key set. Since these two vectors are categorical data, one could use high-order interaction to further explore the useful patterns, but this is out of the scope of this paper. Finally, we feed the embedding of the target sample, retrieved keys, and retrieved values into an MLP to make the label prediction.
@@ 4.5. Time Complexity & Speedup
In this section, we provide a brief time complexity analysis of LIFT framework and then discuss the speedup feasibility. The detailed discussions are deferred to Appendix C.
Time complexity. LIFT consists of two major phases. The first is the pretraining phase targeted at context sequences. In this phase, the dataset $D_{\text{retrieval}}$ is segmented into a set of sequences, each of which is characterized by a length $L$. As the operations on each sequence with a time complexity of $O(1)$, the time complexity of this phase is $O(|D_{\text{retrieval}}|/L)$.
The second phase lies in the LIFT model training and inference. Besides the neural network computation, the retrieval process is incorporated into the model, accounting for a time complexity of $O(F)+O(F\cdot|D_{\text{retrieval}}|/U)=O(F\cdot|D_{\text{retrieval}}|/U)$, where $U$ denotes the total number of unique features encountered in $D_{\text{retrieval}}$, $F$ denotes the number of features in the query.
Speedup. To enhance the efficiency of the retrieval process, we may adopt a method of storing the retrieval outcomes, eliminating the need to conduct the retrieval procedure during online inference. This approach shifts the retrieval process to an offline setting. For instance, in recommender systems, this method involves storing the retrieval results for all (or frequent) users, along with their respective recalled candidate items. This method represents a balance between the complexities of space and time, optimizing resource utilization by trading off storage space for processing speed.
@@ 5. Experiments
This section starts with five research questions (RQs), which we use to guide the experiments and discussions. [^1] We provide the experiment code with running instructions on Anonymous GitHub (https://anonymous.4open.science/r/LIFT-277C/Readme.md).
@@
    How is the time efficiency of LIFT and whether is it potential to be deployed online?
@@ 5.1. Datasets
We evaluate the performance of LIFT by conducting experiments for CTR prediction tasks on three large-scale real-world datasets, i.e., Taobao, Tmall, and Alipay [^2] https://tianchi.aliyun.com/dataset/x, where ‘x’ is ‘649’, ‘42’, and ‘53’ for Taobao, Tmall, and Alipay, respectively.. For top-N ranking, we utilize two widely-used public recommendation datasets: MovieLen [^3] https://grouplens.org/datasets/movielens/1m/ and LastFM [^4] http://ocelma.net/MusicRecommendationDataset/lastfm-1K.html. We count the number of instances, fields of these datasets, which is shown in Table 1.
Table 1. Dataset statistics
Datasets # Instances # Fields Task
Taobao 100,150,807 4 CTR Prediction
Tmall 54,925,331 9 CTR Prediction
Alipay 35,179,371 6 CTR Prediction
Movielens-1M 1,000,209 7 Top-N Ranking
LastFM 18,993,371 5 Top-N Ranking
For dataset preprocessing, we follow the common practice in RIM(Qin et al., 2021) and DERT (Zheng et al., 2023). We split each dataset into three parts, i.e., retrieval set, train set, and test set, based on the global timestamps. We select the hyperparameters using cross validation over the training set. Specifically, the retrieval set consists of the earliest data instances, the test set comprises the latest data instances, and the remaining intermediate data instances are allocated to the train set. For pretraining, we use the retrieval set to pretrain the encoder. For the non-retrieval baseline models, we merge the retrieval set and the training set as the final training set.
@@ 5.2. Evaluation Metrics
We choose the widely used metrics area under the ROC curve (AUC) and negative log-likelihood (LogLoss) to evaluate the performance for CTR prediction. For top-N ranking, we use hit ratio (HR@N), normalized discounted cumulative gains (NDCG@N), and mean reciprocal rank (MRR). Significance test on each metric between the first and second performed methods is conducted, with * marked for positive test results. We also conduct a significance test on the metrics between LIFT and the best performed baseline, marked with “*” for positive results.
@@
Table 3. Performance comparison of top-N ranking tasks on the ML-1M and LastFM datasets. The best results are highlighted in bold, while the second-best results are underlined. Significant improvements are indicated by $p<0.01$.
Datasets Metric FPMC TransRec NARM GRU4Rec SASRec RIM DERT LIFT
ML-1M NDCG@5 0.0788 0.0808 0.0866 0.0872 0.0981 0.1577 0.1634 0.1806
NDCG@10 0.1184 0.1217 0.1254 0.1265 0.1341 0.2059 0.2117 0.2293
MRR 0.1041 0.1078 0.1113 0.1135 0.1193 0.1704 0.1774 0.1914
HR@1 0.0261 0.0275 0.0337 0.0369 0.0392 0.0645 0.0747 0.0800
HR@5 0.1334 0.1375 0.1418 0.1395 0.1588 0.2515 0.2540 0.2808
HR@10 0.2577 0.2659 0.2631 0.2624 0.2709 0.4014 0.4035 0.4324
LastFM NDCG@5 0.0432 0.1148 0.0916 0.1229 0.1163 0.2165 0.2620 0.2723
NDCG@10 0.0685 0.1441 0.1185 0.1486 0.1409 0.2911 0.3217 0.3444
MRR 0.0694 0.1303 0.1083 0.1362 0.1289 0.2210 0.2694 0.2727
HR@1 0.0148 0.0563 0.0423 0.0658 0.0584 0.0915 0.1488 0.1485
HR@5 0.0733 0.1725 0.1394 0.1785 0.1729 0.3468 0.3742 0.4010
HR@10 0.1531 0.2628 0.2227 0.2581 0.2499 0.5780 0.5597 0.6310
@@
Table 5. The influence of the usage of future data in LIFT.
@@ 5.3. Compared Methods
In CTR prediction, we compare LIFT with nine solid baselines that can be categorized into three groups. The first group is traditional tabular models, which do not utilize the sequential or retrieval mechanism. This group includes GBDT (Chen et al., 1996), a widely-used gradient-boosted trees model, and DeepFM (Guo et al., 2017), a factorization-machine based deep neural network. The second group consists of end-to-end sequential deep models, such as DIN (Zhou et al., 2018b) and DIEN (Zhou et al., 2019) that are attention-based recurrent neural networks for CTR prediction, and BERT4Rec (Sun et al., 2019) that is a Transformers-based sequential recommendation model. The third group includes retrieval-based models, such as SIM (Qi et al., 2020), UBR (Qin et al., 2020), RIM (Qin et al., 2021) and DERT(Zheng et al., 2023). Additionally, FATE (Wu et al., 2021) is a tabular data representation learning model that can be viewed as a random retrieval method for prediction.
For top-N ranking, we compare LIFT against seven baseline models. FPMC (Rendle, 2010) and TransRec (He et al., 2017) are factorization-based approaches. The other baselines—NARM (Li et al., 2017), GRU4Rec (Hidasi et al., 2015), SASRec (Khandelwal et al., 2019), RIM (Qin et al., 2021) and DERT (Zheng et al., 2023)—are recently proposed neural network models.
@@ 5.4. Overall Performance (RQ1)
The overall performance comparison result is provided in Table 2, where we have the following observations: (i) LIFT consistently outperforms all nine baselines. Specifically, compared to the best baseline RIM, LIFT achieves a relatively improved AUC by 3.43%, 1.17% and 9.97% on Taobao, Tmall, and Alipay, respectively. This clearly demonstrates the effectiveness of context information learned and retrieved in LIFT framework. (ii) Without pretraining, LIFT still consistently outperforms the best baseline, which indicates that the raw information of future and history still yields a significant improvement on the prediction performance. (iii) The the retrieval-based methods LIFT and RIM are superior to the traditional methods and sequential CTR models, which means that the retrieval methods are able to make better use of the context information. (iv) in the table 3, it can be observed that LIFT achieves significant improvements over these baselines across nearly all metrics on both datasets. This demonstrates that incorporating contextual information enables the sequential model to perform well in top-N ranking tasks.
@@ 5.5. Further Analysis
Future Context (RQ2). Table 5 compares the different parts of the context information’s impact on the final prediction performance. We can observe that both the history information and future information significantly impact on the final prediction. If we only use part of the context information, such as future only or history only, we get a worse result than using both of them, which indicates that the future part of the context provides different information compared with the history part. This observation is entirely neglected by previous works.
Pretraining (RQ3). From the last two rows of Table 2, we find that over the three datasets, the model with pretraining performs better than those without, which clearly demonstrates the efficiency of pretraining. The enhanced performance of the pretrained encoder implies that relying solely on the label provided by the target sample is insufficient for effectively capturing contextual information. To improve the modeling of context, it is imperative to harness the inherent information embedded within the context itself.
@@
The Choice of the Encoder (RQ4). The prediction performance of LIFT with different encoders is provided in Table 4, where we can observe that the Transformer decoder based encoder yields the best performance on both non-pretraining and pretraining settings. The reason could be that the Transformer decoder could be aware of the position of the sequence better. From Figure 4, we observe the mask ratio’s influence on the prediction performance. Usually, the traditional models that encode the user interaction sequence only use the last label as the supervised signal. From the result, we can find that in all three datasets, such a $1$ label supervised strategy is not the best choice. The optimal mask label rate is different over datasets where the number is usually larger than 50%.
Time Efficiency (RQ5). To evaluate the time efficiency of LIFT, we compare its real inference time against other three mainstream models in different types. The parameters that performed best on Alipay were selected for each model to test the inference performance. As shown in Figure 5, LIFT does not introduce significant overload compared with the basic retrieval-based model RIM. The optimal retrieval set size $K$ for LIFT’s AUC on the Alipay dataset is 15. RIM with $K=10$ is approximately 30% faster than LIFT. With offline pretraining, LIFT operates far more efficiently than the transformer-based model BERT4Rec.
Figure 5. Inference time on Alipay.
6. Conclusion
In this work, we propose a retrieval-based framework called LIFT to better utilize the context information of the current user interaction. We are the first work to include future information as part of the context without temporal data leakage in the training and inference stages. Moreover, we use a pretraining method to better mine the information in the context sequences and propose a novel mask behavior loss. The performance of the LIFT framework shows that both historical and future information yield significant improvements in the CTR prediction and top-N ranking performance. Also, from the comparison of the pretraining method, we find that context pretraining is a promising solution to further improve the prediction performance of a variety of sequential recommendation models. In the future, we will investigate deeper on context representation learning with more sophisticated retrieval methods and the speedup schemes to deploy LIFT to real-world recommenders.
Appendix A Notations
The notations and their descriptions are summarized in Table 6.
Table 6. Notations and corresponding descriptions.
Notation Description
$x_{z}$, $\mathbf{x}_{z}$ The raw feature and the embedding of $z$-th sample
$x_{t}$, $\mathbf{x}_{t}$ The raw feature and the embedding of target sample
$E_{\omega}$ The encoder and its parameters
$R$ The retriever
$F_{\theta}$ The predictor and its parameters
$d$ An interaction
$v$ The encoder output vector dimension
$w$ The feature embedding vector dimension
$K$ The size of retrieved contexts
$L$ The length of history and future sequence
$c$, $h$, $f$ The context sequence, history sequence, future sequence
Appendix B Datastore of LIFT
As shown in Figure 6, the datastore $(\mathcal{K},\mathcal{V})$ is the set of all key-value pairs constructed from all the samples in the retrieval dataset $\mathcal{D}_{\text{retrieval}}$:
Figure 6. The datastore in LIFT.
The primary keys within this datastore are intricately linked to the interactions in the dataset $\mathcal{D}_{\text{retrieval}}$, which represents the original data format. In parallel, the datastore houses values that align with the contextual information associated with these interactions. It is noteworthy that this contextual information is derived through an encoding process that applies a pretrained encoder, denoted as $E$, to the context sequences.
Appendix C Time complexity
We aim to provide a comprehensive analysis of the time complexity within our proposed framework, which consists of two major components.
Appendix D Hyperparameter in the Retriever
In the LIFT framework, $L$ means the context sequence length and $K$ means the retrieved samples count. From Figure 7 and Figure 8, we can find that with the increase of $K$, the final prediction result first goes up and then down. The curve indicates that along with with the $K$, the information first increases, and then more noises are introduced to become the dominant part. The AUC curves of $L$ show the similar trends of $K$, i.e., as $L$ increases, the AUC first gets better and then drops down. We think it is the same reason as for $K$. Moreover, because of the limitation of the GPU resources, we only conduct this hyperparameter study on Taobao and Alipay. and we could only increase the $L$ to 70 in the history length of Taobao, where the downtrend just started.